OpenStudy (anonymous):

Given f(x) = x + 2 and g(x) = x2 – 4, find the function (g/f)(x). A. (g/f)(x) = x – 4, excluded value: x = -4. B. (g/f)(x) = x – 2, excluded value: x = -2. C. (g/f)(x) = x – 4, excluded value: x = 4. D. (g/f)(x) = x – 2, excluded value: x = 2.

OpenStudy (anonymous):

okay well we can factor g(x) to be (x+2)*(x-2)

OpenStudy (anonymous):

so (g/f)(x) would be \[\frac{ (x+2)(x-2) }{ (x+2) }\]

OpenStudy (anonymous):

so the (x+2) in the numerator and denominator cancel out, and you're left with x-2

OpenStudy (anonymous):

so (g/f)(x) = x-2

OpenStudy (anonymous):

but x cannot equal -2, because if x = -2, the denominator would be 0, and we cannot divide by 0

OpenStudy (anonymous):

so your answer is B

OpenStudy (anonymous):

Oh wow thank you! :)

OpenStudy (anonymous):

you're welcome